The sum of the ages of two brothers Is 38.four years ago,the age of the elder brother was the square of the younger brother's age.find the product of their ages
Answer by Boreal(15235)

younger=x
older=38-x
;
34-x=(x-4)^2
34-x=x^2-8x+16
0=x^2-7x-18
(x-9)(x+2)=0
x=9; ignore x=-2
The ages are 9 and 29, the product (answer) is 261.
Four years ago, the ages were 5 and 25. Check.